OpenStudy (anonymous):

5(2x-10)-3=5(x-7)+(2)

sam (.sam.):

5(2x-10)-3=5(x-7)+(2) Move all terms not containing x to the right-hand side of the equation. 5(2x-10)=-30+5x Multiply 5 by each term inside the parentheses. 10x-50=-30+5x 5x-50=-30 5x=20 x=4